1. Picasso and Pablo live in Greenwich Village. They are both struggling artists. Picasso is the more productive artist. He can produce 15 small sculptures each day but only 5 paintings. Pablo can produce 4 sculptures each day but only 2 paintings

a) What is the opportunity cost of a painting for each artist?

b) Based on your answer in part (a), who has a comparative advantage?

c) If the two artists were to decide to specialize, who should produce the sculptures and who should produce the painting?

Im having trouble with this practice question for my upcoming final.
Answer #1

Ans.

a- Opportunity cost of a painting for Picasso = 15/5 = 3 small sculpture

Opportunity cost of a painting for Pablo = 4/2 = 2 sculpture

b- Since opportunity cost of producing a painting is lower for Pablo, so Pablo has a comparative advantage in Producing painting.

Picasso has a comparative advantage in Producing sculpture.

c) They will specialize in the good in which they have a comparative advantage.

So, Pablo will specialize in Producing painting.

and Picasso will specialize in producing sculpture.